# ConstScene: A Dataset and Model for Advancing Robust Semantic Segmentation in Construction Environments

### 27 Dec 2023 v1

@@ 2 Related Work
Table 1 compares ConstScene with off-road autonomous driving datasets. Compared to prior studies, ConstScene for the first to address semantic segmentation in challenging environments, such as rain, dust, and fog. Plus, we employ data augmentation techniques to robustify the model training against perturbed data. Notably, our dataset is openly accessible, ensuring result reproducibility. This facilitates researchers to focus on introducing novel detection techniques rather than repeating the time-consuming data collection process.
@@ 3.1 Data Collection and Sensor Setup
The selection of objects in image datasets is vital for training models to precisely recognize elements such as pedestrians, machines, and piles, ultimately enhancing the reliability and decision-making capabilities of self-driving machines in diverse and dynamic construction environments. In this paper, we try to include commonly used objects in construction environments, including construction machines (wheel-loaders), crushers, humans, piles, roads, and background regions.
The image data collection process poses several challenges, including variability in lighting and weather conditions, diverse backgrounds, occlusions, and the need to include authentic situations such as existing dust on the camera lens or workers climbing the ladder of a heavy machine. To address these challenges, we carefully consider the selection of representative samples from diverse environments to ensure model generalization. Plus, we employ data augmentation (Section 3.3), where images are artificially manipulated to simulate variations in weather conditions. Occlusion challenges are mitigated by collecting images from multiple angles, while meticulous labeling and annotation processes are undertaken to provide the necessary ground truth for training robust models. For collecting image data, we utilized an RGB camera featuring a 1280$\times$720 resolution, capturing images in JPEG format at a rate of 15 Frame-Per-Second.
@@ 3.2 Data Annotation
We leverage Roboflow [9] tool to annotate images. The tool allows us to label objects with semantic segmentation masks, enabling precise annotations for object classes. For each image in the dataset, a single Portable Network Graphics (PNG) mask is generated to show the segmented regions corresponding to different classes, providing a representation of the pixel-level annotations. After labeling the whole dataset, images were randomly divided into three categories train, validation, and test.
@@ 3.3 Data Augmentation
Image augmentation plays a crucial role in enhancing the robustness and generalization capabilities of semantic segmentation models [2]. By diversifying the training dataset through changes in contrast, and/or adding perturbations using noise, ultimately we improve the model’s ability to perform well on unseen or slightly different data during testing. In this paper, we leverage three data augmentation techniques including (i) blurring images by taking 25 neighbor pixels and averaging them to perturb images that are not visible to the human eye; (ii) adding pepper-and-salt noise to 1% of pixels; and (iii) removing eight random regions with the size of 3% of image size through a cut-out process for simulating dirty and muddy camera lens. The three first augmentation techniques have been applied to all images as shown in multiple samples in Fig. 1.
@@ 3.4 Depth Estimation
Depth estimation is the task of determining the distance from the camera to objects in a scene [15, 16, 18]. Depth estimation is crucial for various applications, such as robotics, autonomous machines, and augmented/virtual reality. Given the significance of depth estimation in autonomous driving, we provide depth maps of images in the dataset using a monocular depth estimation method [20]. Fig. 2 shows multiple RGB samples of ConstScene with corresponding depth maps across diverse weather and working conditions. The red circles in Fig. 2.c to Fig. 2.e mark the areas of disparity with low-confidence matches for conditions with the presence of fog, dust, and dirt on the camera lens. Consequently, the impact of adverse weather conditions significantly diminishes the accuracy of depth estimation, emphasizing the crucial need to address this challenge in computer vision pipelines employed in construction environments.
@@ 5.3 Prediction Results
Analyzing Results of U-Net. Table 3 presents the test results of U-Net with various backbone architectures trained on the original ConstScene with and without the incorporation of data augmentation techniques. Results show that U-Net with the ResNet50 encoder trained on the non-augmented dataset ($D1$) provides the best results (55.48%) compared to other training settings. In addition, data augmentation ($D2$) provides up to 11.59% higher accuracy for U-Net with the ResNet50 encoder demonstrating the model better learns variations present in real-world data.
@@
Analyzing Results of SegFormer. Table 4 shows the test results of SegFormer trained on ConstScene. The findings reveal that SegFormer-B5, when trained on the original dataset ($D1$), achieves an accuracy of 82.27%, surpassing SegFormer-B5’s results by an improvement of 35.3%.
Analyzing Results of Data Augmentation. When SegFormer-B5 undergoes training using the dataset enhanced with data augmentation $D2$, it yields superior accuracy compared to the non-augmented dataset. This suggests that the $blurring+pepper-and-salt~{}noise+regions~{}removal$ augmentation technique contributes to the model’s improved ability to learn variations present in real-world data.
@@ 6 Conclusion and Future Work
We presented a novel large-scale dataset, dubbed ConstScene, for studying semantic object detection in construction environments across various weather and environmental conditions with extensive annotations for comprehensive understanding and analysis. We extend our dataset by employing data augmentation techniques aligned with environmental hazards prevalent in construction sites. We also showed the results of two popular semantic segmentation models trained on our proposed dataset. In future work, semantic segmentation in construction environments can be further improved by (i) considering objects from other construction segments such as mining; and (ii) generating synthetic data using generative adversarial networks. We hope our new findings and the proposed dataset will stimulate the development of innovative techniques to enhance the efficacy of object detection models against adversarial attacks.

### 19 Jan 2024 v2

@@ 2 Related Work
Table 1 compares ConstScene with off-road autonomous driving datasets. Compared to prior studies, ConstScene is the first to address semantic segmentation in challenging environments, such as rainy, dusty, and fogy. Plus, we employ data augmentation techniques to robustify the model training against perturbed data. Notably, our dataset is openly accessible, ensuring result reproducibility. This facilitates researchers to focus on introducing novel detection techniques rather than repeating the time-consuming data collection process.
@@ 3.1 Data Collection and Sensor Setup
The selection of objects in image datasets is vital for training models to precisely recognize elements such as human, machines, and piles, ultimately enhancing the reliability and decision-making capabilities of self-driving machines in diverse and dynamic construction environments. In this paper, we try to include commonly used objects in construction environments, including construction machine (wheel-loader), crusher, human, pile, road, and background regions.
The image data collection process poses several challenges, to include variability in the dataset such as lighting and weather conditions, diverse backgrounds, occlusions, and the need to include authentic situations such as existing dust on the camera lens or workers climbing the ladder of a heavy machine. To address these challenges, we carefully have selected representative samples considering diverse environment conditions to ensure model generalization. Plus, we employ data augmentation (Section 3.3), to artificially manipulate the input images to simulate variations in weather conditions. Occlusion challenges are mitigated by collecting images from multiple angles, while meticulous labeling and annotation processes are undertaken to provide the necessary ground truth for training robust models. For collecting image data, we utilized an RGB camera featuring a 1280$\times$720 resolution, capturing images in JPEG format at a rate of 15 Frame-Per-Second.
@@ 3.2 Data Annotation
We leverage Roboflow [9] tool to annotate images. The tool allows us to label objects with semantic segmentation masks, enabling precise annotations for object classes. For each image in the dataset, a single Portable Network Graphics (PNG) mask is generated to show the segmented regions corresponding to different classes, providing a representation of the pixel-level annotation. After labeling the whole dataset, images were randomly divided into three categories train, validation, and test.
@@ 3.3 Data Augmentation
Image augmentation plays a crucial role in enhancing the robustness and generalization capabilities of semantic segmentation models [2]. By diversifying the training dataset through changes in contrast, and/or adding perturbations using noise, ultimately we improve the model’s ability to perform well on unseen or slightly different data during testing. In this paper, we leverage three data augmentation techniques by (i) blurring images through taking 25 neighbor pixels and averaging them; (ii) adding pepper-and-salt noise to 1% of pixels; and (iii) removing eight random regions with the size of 3% of image size through a cut-out process for imitating dirty and muddy camera lens. In the first two methods the added perturbation may not be visible to human eye but still can make the model to miss-classify. These augmentation techniques have been applied to the images of the train set in the dataset. Fig. 1 shows several samples from the dataset where these three perturbation techniques are applied.
@@ 3.4 Depth Estimation
Depth estimation is the task of determining the distance from the camera to objects in a scene [15, 16, 18]. Depth estimation is crucial for various applications, such as robotics, autonomous machines, and augmented/virtual reality where environment perception is required for decision making. In autonomous machines application, depth estimation plays a crucial role for obstacle avoidance that contributes to safety. Given the significance of depth estimation in autonomous driving, we provide depth maps of images in the dataset using a monocular depth estimation method [20]. Fig. 2 shows multiple RGB samples of ConstScene with corresponding depth maps across diverse weather and working conditions. The red circles in Fig. 2.c to Fig. 2.e mark the areas of disparity with low-confidence matches for conditions with the presence of fog, dust, and dirt on the camera lens. As can be seen, the impact of adverse weather conditions significantly diminishes the accuracy of depth estimation, emphasizing the crucial need to address this challenge in computer vision pipelines employed in construction environments.
@@ 5.3 Prediction Results
Analyzing Results of U-Net. Table 3 presents the test results of U-Net with various backbone architectures trained on the original ConstScene with and without the incorporation of data augmentation techniques. Results show that U-Net with the ResNet50 encoder trained on the non-augmented dataset ($D1$) provides the best results (55.48%) compared to other training settings. In addition, data augmentation ($D2$) provides up to 11.59% higher accuracy for U-Net with the ResNet50 encoder demonstrating that the model excels at capturing the diverse patterns found in real-world.
@@
Analyzing Results of SegFormer. Table 4 shows the test results of SegFormer trained on ConstScene. The findings reveal that SegFormer-B5, when trained on the original dataset ($D1$), achieves an accuracy of 82.27%, surpassing SegFormer-B0’s results by an improvement of 35.3%.
Analyzing Results of Data Augmentation. When SegFormer-B5 undergoes training using the dataset enhanced with data augmentation $D2$, it yields superior accuracy compared to the non-augmented dataset. This suggests that the $blurring+pepper-and-salt~{}noise+regions~{}removal$ augmentation techniques contribute to the model’s improved ability to learn variations present in real-world data.
@@ 6 Conclusion and Future Work
We presented a novel large-scale dataset, dubbed ConstScene, for studying semantic object detection in construction environment across various weather and environmental conditions with extensive annotations for comprehensive understanding and analysis. We have extended our dataset by employing data augmentation techniques aligned with environmental hazards prevalent in construction sites. We also showed the results of two popular semantic segmentation models trained on our proposed dataset. In future work, semantic segmentation in construction environments can be further improved by (i) considering objects from other construction segments such as mining; and (ii) generating synthetic data using generative adversarial networks. Our new findings and the proposed dataset pose a high potential in stimulating the development of innovative techniques to enhance the efficacy of object detection models against adversarial attacks.